A network is given as 190.28.0.0/16. Eight subnets are to be created.

(a)What will be the CIDR network prefix?

(b)Express last subnet in dotted decimal notation with mask.

(c) What is the maximum number of hosts in each subnet?

(d) What is the address range for hosts in second subnet?

(e) What is the subnet broadcast address for the first subnet?


1
Expert's answer
2022-03-29T11:40:57-0400

190.28.0.0/16

 

/16 means that Subnet Mask:           255.255.0.0


Total number of host is in range of -.-.xxx.yyy mask, where xxx can be 0-255 (256-1 “positions”), yyy also can have value 0-255 (256-1 “positions”).

xxx = 0, it means that have maximum 256 “positions”

yyy = 0, it means that have maximum 256 “positions”

Total Number of Hosts = 256 x 256 = 65 536

Network with “/z” prefix (subnet mask) has number of hosts = 2(32 – z)

32 – 16 = 16

216 = 65 536

Other simple example:

Mask 255.255.255.240

We can see that xxx have maximum value 255 (minimum “position”), so, number of host should be determinate in yyy range.

yyy = 240, total number of host will be 256 – 240 = 16

 

Eight subnet should be created

65536 : 8 = 8192, so in each subnet’s total number of host will be 8192.

We can see that it is more than 256, so yyy = 0 (have maximum 256 “positions”)

How many 256 from 8192 we have, i.e. how xxx value we have which in difference with 256 will provide 256 count in 8192?

8192 : 256 = 32, so, xxx can have value = 256 – 32 = 224.

Our subnet with total number of host = 8192 will have subnet mask = 255.255.224.0

Network with “/z” prefix (subnet mask) has number of hosts = 2(32 – z)

2(32 – z) = 8192

32 - z = 13

z = 19

Subnet mask = 255.255.224.0 means that we have “/19” prefix.

Total Number of Hosts:        8,192

Number of Usable Hosts:      8,190

Subnet Mask:                        255.255.224.0

CIDR Notation:                    /19


All 8 of the Possible /19 Networks for 190.28.*.*

Network Address/Usable Host Range/Broadcast Address:

1) 190.28.0.0/190.28.0.1 - 190.28.31.254/190.28.31.255

2) 190.28.32.0/190.28.32.1 - 190.28.63.254/190.28.63.255

3) 190.28.64.0/190.28.64.1 - 190.28.95.254/190.28.95.255

4) 190.28.96.0/190.28.96.1 - 190.28.127.254/190.28.127.255

5) 190.28.128.0/190.28.128.1 - 190.28.159.254/190.28.159.255

6) 190.28.160.0/190.28.160.1 - 190.28.191.254/190.28.191.255

7) 190.28.192.0/190.28.192.1 - 190.28.223.254/190.28.223.255

8) 190.28.224.0/190.28.224.1 - 190.28.255.254/190.28.255.255

 

Full info of first subnet

1)     190.28.0.0/190.28.0.1 - 190.28.31.254/190.28.31.255

Network Address:                190.28.0.0

Usable Host IP Range:          190.28.0.1 - 190.28.31.254

Broadcast Address:               190.28.31.255

Total Number of Hosts:        8,192

Number of Usable Hosts:      8,190

Subnet Mask:                       255.255.224.0

CIDR Notation:                    /19

Short:                                    190.28.0.0 /19

 

(a)What will be the CIDR network prefix?

/19

(b)Express last subnet in dotted decimal notation with mask.

8) 190.28.224.0/190.28.224.1 - 190.28.255.254/190.28.255.255

Network Address:                190.28.224.0

Usable Host IP Range:          190.28.224.1 - 190.28.255.254

Broadcast Address:               190.28.255.255

Total Number of Hosts:        8,192

Number of Usable Hosts:      8,190

Subnet Mask:                        255.255.224.0

CIDR Notation:                    /19

Short:                                    190.28.224.0 /19

(c) What is the maximum number of hosts in each subnet?

Total Number of Hosts:        8,192

Number of Usable Hosts:      8,190

 

(d) What is the address range for hosts in second subnet?

2) 190.28.32.0/190.28.32.1 - 190.28.63.254/190.28.63.255

Usable Host IP Range: 190.28.32.1 - 190.28.63.254

(e) What is the subnet broadcast address for the first subnet?

1) 190.28.0.0/190.28.0.1 - 190.28.31.254/190.28.31.255

190.28.31.255